MILLIMETER WAVE SPECTROSCOPY OF HIGH RYDBERG STATES OF Kr

\underline{MARTIN SCHÄFER}, GREGORY M. GREETHAM, FRÉDÉRIC MERKT, ETH Zürich, Laboratorium für Physikalische Chemie, ETH-Hönggerberg HCI, CH-8093 Zürich, Switzerland.

A phase-stabilized backward wave oscillator (BWO) in the 260–380 GHz range was combined with an XUV laser system to record high-resolution spectra of high Rydberg states of krypton. The millimeter wave transitions between high n Rydberg states were detected by pulsed field-ionization, at sub-MHz resolution. Results are presented on the hyperfine structure of high n Rydberg states of ^{83}Kr.